# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ 1.2. Our contributions
We give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner (HSS) [13]. This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantiations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms). Combining our concurrent composition theorems (both Theorem 1.2 and Corollary 1.7), we obtain a much simpler privacy proof. Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
Our Variants of Concurrent Composition.
Here we elaborate on the different types of concurrent composition that we study in this paper, with more general and more precise formulations than described above.
@@
The paper is organized as follows. Section 2 introduces all necessary notation, including non-interactive mechanisms, interactive mechanisms, and interactive post-processing mechanism. Section 3 defines continual mechanisms, identifiers $\mathcal{I}$, verification functions $f$, verifiers $\mathcal{V}$ and differential privacy for continual mechanisms. Concurrent composition of continual mechanisms is defined in Section 4 and our first concurrent composition theorem is proven. Section 5 introduces concurrent $k$-sparse parallel composition of continual mechanisms and interactive mechanisms, proves the theorems for this type of composition and shows the counterexample. In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use non-interactive, interactive and continual sub-mechanisms. In Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of [13]. Finally, in Section 8 and Section 9, we prove two technical intermediate lemmas, required for the previous results.
@@ 2.1. Non-interactive Mechanisms (NIM)
Definition 2.1 ($(\epsilon,\delta)$-indistinguishability [17]).
For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, two random variables $X_{1}$ and $X_{2}$ over the same domain $\mathcal{Y}$ are $(\epsilon,\delta)$-indistinguishable if for any subset $Y\subseteq\mathcal{Y}$,
@@
Definition 2.2 ($(\epsilon,\delta)$-DP for NIMs [6]).
@@
The reason [16] introduced $\mathrm{RR}_{\epsilon,\delta}$ is that it is ”universal” among $(\epsilon,\delta)$-DP NIMs in the following sense:
Theorem 2.5 ([16]).
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP w.r.t. a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized ”post-processor” $\mathcal{P}:\mathcal{Y}\to\{0,1\}\times\{\top,\bot\}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta}(b))$.
@@ 2.2. Interactive Mechanisms (IM)
In the differential privacy literature, a query typically refers to a question about the dataset held by a mechanism. However, in this paper, we use query as a shorthand for query message, meaning an input message to a mechanism that requires a response.
@@
An example of an IM is the interactive randomized response mechanism, denoted $\mathrm{IRR}_{\epsilon,\delta}$, which will repeatedly be used in our proofs. Ghazi et al. [10] define this mechanism to initially flip a coin and determine whether to answer all subsequent queries truthfully or in a privacy-preserving manner. Our definition of interactive randomized response mechanism is different but shares the same intuition.
@@
Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be two interactive mechanisms with $Q_{\mathcal{M}}=Q_{\mathcal{M}^{\prime}}$ and $A_{\mathcal{M}}=A_{\mathcal{M}^{\prime}}$. Let $s$ and $s^{\prime}$ be initial states for $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively. The mechanisms $\mathcal{M}(s)$ and $\mathcal{M}^{\prime}(s^{\prime})$ are equivalent if for any adversary $\mathcal{A}$ with $A_{\mathcal{A}}=Q_{\mathcal{M}}$ and $Q_{\mathcal{A}}=A_{\mathcal{M}}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}^{\prime}(s^{\prime}))$ are identically distributed.
@@
Existing works [23, 24, 19, 11] view the initial state of $\mathcal{M}$ is a dataset, i.e., $S_{\mathcal{M}}^{\text{init}}=\mathcal{X}$, and (as is usual) they define neighbor relations as binary relations on $\mathcal{X}$. However, for us it will be convenient to conceptually separate $S_{\mathcal{M}}^{\text{init}}$ from the data (which is changing continuously over time), but still define DP w.r.t. neighbor relations on $S_{\mathcal{M}}^{\text{init}}$. This leads to the following definition of differential privacy for IMs.
@@ 2.3. Interactive Post-Processing (IPM)
Haney et al. [11] generalize post-processing functions to interactive post-processing mechanisms (IPMs) as follows and show that interactive post-processing preserves privacy guarantees for IMs.
Definition 2.14 (Interactive Post-Processing Mechanism (IPM) [11]).
@@
Definition 2.15 (Post-Processing of an IM $\mathcal{M}$ by an IPM $\mathcal{P}$ ($\mathcal{P}\!\circ^{*}\!\mathcal{M}$)[11]).
@@
Algorithm 1 Post-Processing of $\mathcal{M}$ by $\mathcal{P}$, denoted by $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ [11].
@@
Haney et al. [11] shows that if an IM $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private, then for any IPM $\mathcal{P}$ with a singleton initial state space, $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-differentially private. This result is formally stated below.
Lemma 2.16 (Post-Processing Lemma for IMs [11]).
Let $\mathcal{M}$ be an IM $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$. For every pair of initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$, the following holds: If for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable, then for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable.  [^5] In [11], the definitions of IMs and IPMs are more restrictive. Specifically, for any IM $\mathcal{M}$, $S_{\mathcal{M}}=Q_{\mathcal{M}}=A_{\mathcal{M}}=\{0,1\}^{*}$, and for any IPM $\mathcal{P}$, $S_{\mathcal{P}}=M_{\mathcal{P}}=\{0,1\}^{*}$; However, their results do not rely on these restrictions, and the interactive post-processing lemma holds for arbitrary sets $S_{\mathcal{M}}$, $Q_{\mathcal{M}}$, $A_{\mathcal{M}}$, $S_{\mathcal{P}}$, and $M_{\mathcal{P}}$.
@@ 3. Continual Mechanisms
This section is based on two crucial observations: (i) A message, as sent or received by IMs and IPMs, is a general and powerful concept capable of encoding various types of instructions, and (ii) the essence of DP is to protect the privacy of a single bit that distinguishes neighboring states, sequences, etc. The second observation has been used many times in the DP literature and was specifically used in [14] to model adaptivity in continual observation. Using these key observations, we first reformulate the definition of DP for IMs and continual observation. Then, inspired by the new formulations, we introduce continual mechanisms (CMs) and formally define DP in this setting.
@@
By Definition 2.13, an interactive mechanism $\mathcal{M}$ with initial state space $S_{\mathcal{M}}^{\text{init}}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$, if for every two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and every adversary $\mathcal{A}$, the view of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ is $(\epsilon,\delta)$-indistinguishable. This means that for the worst-case (i.e., adversarial) choice of neighboring initial states, an adversary by adaptively asking queries should not be able to determine whether $\mathcal{M}$ is running on $s_{0}$ or $s_{1}$.
@@
To formalize this, we restrict attention to adversaries that send a pair of neighboring initial states $(s_{0},s_{1})$ as their first message and pairs of identical queries $(q_{i},q_{i})$ in its subsequent messages. We also define an IPM $\mathcal{I}$ with the initial state space $S_{\mathcal{I}}^{\text{init}}=\{0,1\}$ to stand between the adversary and $\mathcal{M}$, filtering the pairs based on its, i.e., $\mathcal{I}$’s, initial state. Specifically, for $b\in\{0,1\}$, in an interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$. With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for every adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
Actually, we will enforce the restrictions on the adversary also by an IPM, which we call a verifier $\mathcal{V}$. The verifier checks whether the initial message is a pair of neighboring datasets and the subsequent messages of the adversary are of the form $(q,q)$; only if these checks pass, are the messages forwarded to $\mathcal{I}$.
@@
Dwork, Naor, Pitassi, and Rothblum [7] introduced differential privacy under continual observation, where the dataset is not given at the initialization time and instead evolves over time. In this model, the mechanism receives either an empty record or a data record at each step. If the record is non-empty, the mechanism adds it to its dataset. Then, whether the input record is empty or not, the mechanism answers a predetermined question about its current dataset.
@@
To capture ”event-level privacy” (where the privacy unit is a single record), Dwork et al. [7] define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in a single step. They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for every two neighboring input sequences. This definition does not allow for data sequences that are adaptively chosen based on the answers released.
@@
Jain, Raskhodnikova, Sivakumar, and Smith [14] formalized the definition of DP against adaptive adversaries in the continual observation model as follows: An adaptive adversary $\mathcal{A}$ chooses a single (data or empty) record at every step, except for one challenge step, selected by the adversary, where $\mathcal{A}$ sends a pair of records. They called a mechanism $\mathcal{M}$ $(\epsilon,\delta)$-differentially private against adaptive adversaries in the continual observation model if for any such adaptive adversary $\mathcal{A}$, the distributions of $\mathcal{M}$’s answer sequences are $(\epsilon,\delta)$-indistinguishable when the first or the second element of the pair in the challenge step is handed to $\mathcal{M}$.
@@
A mechanism $\mathcal{M}$ in the continual observation model can be viewed as an IM with a single possible initial state, representing an empty dataset. The input message set of $\mathcal{M}$ consists of the empty record and all possible data records. Instead of having a challenge step, we can alternatively restrict to adversaries that send pairs of identical records at every step except for one. Under this formulation, the definition of differential privacy proposed by [14] can be restated as follows: A mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP differential privacy against adaptive adversaries if for any such adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously. Again, below we will enforce the restriction on the adversary via another IPM $\mathcal{V}$.
@@
In this alternative definition, all adversary messages are pairs of identical data or empty records, except for a single message that is a pair of different records. Recall that in the alternative definition of DP for IMs, the first message sent by an adversary must be a pair of neighboring initial datasets, while all subsequent adversary messages must be pairs of identical queries. Next, we generalize the definition of differential privacy by allowing arbitrary conditions on the sequence of pairs sent by adversaries, without restricting the content of the messages.
@@
  • Every non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ as a response. Upon receiving any subsequent input message, the CM representing $\mathcal{N}$ halts the communication.
@@
  • Mechanisms such as the Sparse Vector Technique (SVT) [8, 12, 18] and the continual counter [6, 2, 3, 5] in the streaming settings can also be represented as CMs receiving two types of messages: question queries and data-update queries. Question queries request information about the current dataset, while data-update queries consist of instructions for modifying the dataset.

### v4

@@ 1. Introduction
We give a new privacy proof for the continual monotone histogram mechanism of Henzinger, Sricharan, and Steiner (HSS) (henzingerSS2023, ). This mechanism involves concurrent executions of a simpler continual histogram mechanism and multiple instantiations of the Sparse Vector Technique mechanism, and the original privacy analysis was rather intricate even for oblivious adversaries (because the mechanism itself is adaptive in its queries and dataset updates to the building-block mechanisms). Combining our concurrent composition theorems (both Theorem 1.2 and Corollary 1.7), we obtain a much simpler privacy proof. Moreover, for the case of approximate DP, the original proof was only for oblivious adversaries, whereas our proof also applies to adaptive adversaries.
An Application in Local Differential Privacy (LDP)
Consider an untrusted server that concurrently interacts with multiple users, each holding a private dataset, with the goal of computing a function on the users’ aggregated data. A local protocol specifies the interaction rules between the server and the users, where both the server and the users are modeled as interactive mechanisms. Differential privacy for a local protocol is defined in terms of the distribution of the server’s view, consisting of its internal randomness and the transcript of all exchanged messages. Specifically, a protocol $P$ is said to satisfy $(\epsilon,\delta)$-LDP if the server’s view is $(\epsilon,\delta)$-indistinguishable for every pair of neighboring sequences of user datasets.
A useful notion of neighboring in local differential privacy (LDP) considers two assignments of datasets to users that are identical for all but at most $k$ users, where the exceptional users datasets are neighboring with respect to a given binary relation on user datasets. An example of this is a setting in which each user represents a hospital and each patient may have records in at most $k$ hospitals. Analyzing the privacy of local protocols in which the server interacts concurrently with users over multiple rounds can be challenging. We observe that such interactions can be modeled as a concurrent parallel composition of interactive mechanisms, and our new privacy results for this concurrent composition directly yield a simplified privacy analysis for local protocols:
Theorem 1.9 (informal statement).
Let $P$ be a local protocol with $n$ users, and let $\sim^{*}$ be a neighbor relation on (individual) user datasets. Define two sequences of $n$ datasets to be neighbors if they are identical for all but at most $k\in\mathbb{N}$ users, where each differing user’s datasets are neighbors with respect to $\sim^{*}$. Suppose the interactive mechanism corresponding to every user is $(\epsilon_{0},\delta_{0})$-DP with respect to $\sim^{*}$ against adaptive adversaries. If the composition of $k$ noninteractive $(\epsilon_{0},\delta_{0})$-DP mechanisms is $(\epsilon,\delta)$-DP, then the protocol $P$ is $(\epsilon,\delta)$-LDP.
In the formal version of Theorem 1.9 (see Corollary 8.8), we incorporate the fact that the server in $P$ restricts the queries sent to each user and enforces that they satisfy certain properties. Consequently, in Theorem 1.9, each user’s privacy guarantee only needs to hold for adversaries whose queries obey these properties. Furthermore, we provide a more general statement of Theorem 1.9 for an arbitrary choice of neighbor relation on sequences of user datasets in Theorem 8.7. The concurrent composition theorem used in the general result must be selected based on the given neighbor relation.
We illustrate an application of Theorem 1.9 by simplifying the privacy analysis of the local protocol proposed by Dhulipala et al. (dhulipala2025near, ) for the core decomposition problem. In that work, a sequence of datasets $x=(x_{1},\dots,x_{n})$ represents a graph with $n$ vertices, where $x_{i}$ denotes the set of vertices adjacent to vertex $i$. (Users are vertices.) Differential privacy is defined with respect to edge-neighboring: two dataset sequences $x$ and $x^{\prime}$ are considered neighbors if their corresponding graphs differ by a single edge. The edge-neighboring relation can be generalized to a relation $\sim$, where $x\sim x^{\prime}$ if and only if the adjacency sets $x_{i}$ and $x_{i}^{\prime}$ are identical for all vertices $i$ but at most $2$ of them, and for those vertices, $x_{i}$ and $x_{i}^{\prime}$ are add-remove neighbors, meaning that they differ in the presence or absence of a single (adjacent) vertex. By defining $\sim^{*}$ in Theorem 1.9 to be the add-remove neighbor relation and setting $k=2$, we obtain a simpler privacy analysis for the protocol in (dhulipala2025near, ).
We note that Theorem 1.9 and its generalizations are stated for protocols in which each user holds a static dataset. In many real-world settings, however, users may join or leave, and their data may evolve over time. In such cases, users must be modeled more generally as continual mechanisms rather than interactive mechanisms. The privacy guarantees of these protocols can still be computed using our concurrent composition theorems for continual mechanisms.
We also note that in the above definition, all components (i.e., the server and the users) are assumed to be honest, meaning that they follow the protocol, and non-colluding, meaning that they do not share their internal states outside the protocol. In the presence of possible dishonesty or collusion, all users except the target users who hold the private data, together with the server, must be treated as a single adversary that interacts concurrently with the target users. In this case, privacy of the local protocol follows from the concurrent composition of the interactive and continual mechanisms, depending on whether the datasets evolve over time.
@@
The paper is organized as follows. Section 2 introduces all necessary notation, including noninteractive mechanisms, interactive mechanisms, and interactive post-processing mechanism. Section 3 defines continual mechanisms, identifiers $\mathcal{I}$, verification functions $f$, verifiers $\mathcal{V}$ and differential privacy for continual mechanisms. Concurrent composition of continual mechanisms is defined in Section 4 and our first concurrent composition theorem is proven. Section 5 introduces concurrent $k$-sparse parallel composition of continual mechanisms and interactive mechanisms, proves the theorems for this type of composition and shows the counterexample. In Section 6, we state the conditions and the theorem for analyzing the privacy of complex continual mechanisms that use noninteractive, interactive and continual sub-mechanisms. In Section 7, we apply our theorem to analyze the privacy of the continual histogram mechanism of (henzingerSS2023, ). In Section 8, we show how local differential privacy can be analyzed using our concurrent composition theorems. In Section 9, we simplify the privacy analysis of the core decomposition local protocol by (dhulipala2025near, ). In Section 10 and Section 11, we prove two technical intermediate lemmas, required for the previous results. In Section 12, we introduce concurrent filter composition of continual mechanisms and reduce it to the filter composition for noninteractive mechanisms. Finally, in Section 13 and Section 14, we provide analogous results for Rényi DP and $f$-DP, respectively.
@@ 2.1. Non-interactive Mechanisms (NIM)
Definition 2.1 ($(\epsilon,\delta)$-indistinguishability (kasiviswanathan2014semantics, )).
For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, two random variables $X_{1}$ and $X_{2}$ over the same domain $\mathcal{Y}$ are $(\epsilon,\delta)$-indistinguishable if for every measurable subset $Y\subseteq\mathcal{Y}$,
@@
Definition 2.2 ($(\epsilon,\delta)$-DP for NIMs (dwork2006calibrating, )).
@@
The reason (KairouzOV15, ) introduced $\mathrm{RR}_{\epsilon,\delta}$ is that it is ”universal” among $(\epsilon,\delta)$-DP NIMs in the following sense:
Theorem 2.5 ((KairouzOV15, )).
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP w.r.t. a neighbor relation $\sim$ if and only if for any $x_{0}\sim x_{1}$, there exists a randomized “post-processor” $\mathcal{P}:\mathcal{Y}\to\{0,1\}\times\{\top,\bot\}$ such that for both $b\in\{0,1\}$, $\mathcal{M}(x_{b})$ is identically distributed to $\mathcal{P}(\mathrm{RR}_{\epsilon,\delta}(b))$.
@@ 2.2. Interactive Mechanisms (IM)
In the differential privacy literature, a query typically refers to a question about the dataset held by a mechanism. However, in this paper, we use query as a shorthand for query message, meaning any input message to a mechanism that requires a response.
@@
An example of an IM is the interactive randomized response mechanism, denoted $\mathrm{IRR}_{\epsilon,\delta}$, which will repeatedly be used in our proofs. Ghazi et al. (ghazi2025privacy, ) define this mechanism to initially flip a coin and determine whether to answer all subsequent queries truthfully or in a privacy-preserving manner. Our definition of the interactive randomized response mechanism is different but shares the same intuition.
@@
Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be two interactive mechanisms with $Q_{\mathcal{M}}=Q_{\mathcal{M}^{\prime}}$ and $A_{\mathcal{M}}=A_{\mathcal{M}^{\prime}}$. Let $s$ and $s^{\prime}$ be initial states for $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively. The mechanisms $\mathcal{M}(s)$ and $\mathcal{M}^{\prime}(s^{\prime})$ are equivalent if for every adversary $\mathcal{A}$ with $A_{\mathcal{A}}=Q_{\mathcal{M}}$ and $Q_{\mathcal{A}}=A_{\mathcal{M}}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}^{\prime}(s^{\prime}))$ are identically distributed.
@@
Existing works (vadhan2021concurrent, ; vadhan2022concurrent, ; lyu2022composition, ; haney2023concurrent, ) view the initial state of $\mathcal{M}$ as a dataset, i.e., $S_{\mathcal{M}}^{\text{init}}=\mathcal{X}$, and (as is usual) they define neighbor relations as binary relations on $\mathcal{X}$. However, for us it will be convenient to conceptually separate $S_{\mathcal{M}}^{\text{init}}$ from the data (which is changing continuously over time), but still define DP w.r.t. neighbor relations on $S_{\mathcal{M}}^{\text{init}}$. This leads to the following definition of differential privacy for IMs.
@@ 2.3. Interactive Post-Processing (IPM)
Haney et al. (haney2023concurrent, ) generalize post-processing functions to interactive post-processing mechanisms (IPMs) as follows and show that interactive post-processing preserves privacy guarantees for IMs.
Definition 2.14 (Interactive Post-Processing Mechanism (IPM) (haney2023concurrent, )).
@@
Definition 2.15 (Post-Processing of an IM $\mathcal{M}$ by an IPM $\mathcal{P}$ ($\mathcal{P}\!\circ^{*}\!\mathcal{M}$)(haney2023concurrent, )).
@@
Algorithm 1 Post-Processing of $\mathcal{M}$ by $\mathcal{P}$, denoted by $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ (haney2023concurrent, ).
@@
Haney et al. (haney2023concurrent, ) shows that if an IM $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private, then for any IPM $\mathcal{P}$ with a singleton initial state space, $\mathcal{P}\!\circ^{*}\!\mathcal{M}$ is also $(\epsilon,\delta)$-differentially private. This result is formally stated below.
Lemma 2.16 (Post-Processing Lemma for IMs (haney2023concurrent, )).
Let $\mathcal{M}$ be an IM with an initial state space $S_{\mathcal{M}}^{\text{init}}$ and $\mathcal{P}$ be an IPM with a singleton initial state space $S_{\mathcal{P}}^{\text{init}}=\{s\}$. For every pair of initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$, the following holds: If for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable, then for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{0}))$ and $\mathrm{View}(\mathcal{A},\mathcal{P}\!\circ^{*}\!\mathcal{M}(s_{1}))$ are $(\epsilon,\delta)$-indistinguishable.  [^10] In (haney2023concurrent, ), the definitions of IMs and IPMs are more restrictive. Specifically, for any IM $\mathcal{M}$, $S_{\mathcal{M}}=Q_{\mathcal{M}}=A_{\mathcal{M}}=\{0,1\}^{*}$, and for any IPM $\mathcal{P}$, $S_{\mathcal{P}}=M_{\mathcal{P}}=\{0,1\}^{*}$; however, their results do not rely on these restrictions, and the interactive post-processing lemma holds for arbitrary sets $S_{\mathcal{M}}$, $Q_{\mathcal{M}}$, $A_{\mathcal{M}}$, $S_{\mathcal{P}}$, and $M_{\mathcal{P}}$, which are not necessarily countable.
@@ 3. Continual Mechanisms
This section is based on two crucial observations: (i) A message, as sent or received by IMs and IPMs, is a general and powerful concept capable of encoding various types of instructions, and (ii) the essence of DP is to protect the privacy of a single bit that distinguishes neighboring states, sequences, etc. The second observation has been used many times in the DP literature and was specifically used in (pricedpco, ) to model adaptivity in continual observation. Using these key observations, we first reformulate the definition of DP for IMs and continual observation. Then, inspired by the new formulations, we introduce continual mechanisms (CMs) and formally define DP in this setting.
@@
By Definition 2.13, an interactive mechanism $\mathcal{M}$ with initial state space $S_{\mathcal{M}}^{\text{init}}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$, if for every two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and every adversary $\mathcal{A}$, the views of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ are $(\epsilon,\delta)$-indistinguishable. This means that for the worst-case (i.e., adversarial) choice of neighboring initial states, an adversary by adaptively asking queries should not be able to determine whether $\mathcal{M}$ is running on $s_{0}$ or $s_{1}$.
@@
To formalize this, we restrict attention to adversaries that send a pair of neighboring initial states $(s_{0},s_{1})$ as their first message and pairs of identical queries $(q_{i},q_{i})$ in its subsequent messages. We also define an IPM $\mathcal{I}$ with the initial state space $S_{\mathcal{I}}^{\text{init}}=\{0,1\}$ to stand between the adversary and $\mathcal{M}$, filtering the pairs based on its, i.e., $\mathcal{I}$’s, initial state. Specifically, for $b\in\{0,1\}$, in the interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$. With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for every adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
We will enforce the restrictions on the adversary by an IPM, which we call a verifier $\mathcal{V}$. The verifier checks whether the initial message is a pair of neighboring datasets and the subsequent messages of the adversary are of the form $(q,q)$; only if these checks pass are the messages forwarded to $\mathcal{I}$.
@@
Dwork, Naor, Pitassi, and Rothblum (Dwork2010, ) introduced differential privacy under continual observation, where the dataset is not given at the initialization time and instead evolves over time. In this model, the mechanism receives either an empty record or a data record at each step. If the record is non-empty, the mechanism adds it to its dataset. Then, whether the input record is empty or not, the mechanism answers a predetermined question about its current dataset.
@@
To capture “event-level privacy” (where the privacy unit is a single record), Dwork et al. (Dwork2010, ) define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in at most one step. They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for every two neighboring input sequences. This definition does not allow for data sequences that are adaptively chosen based on the answers released.
@@
Jain, Raskhodnikova, Sivakumar, and Smith (pricedpco, ) formalized the definition of DP against adaptive adversaries in the continual observation model as follows: An adaptive adversary $\mathcal{A}$ chooses a single (data or empty) record at every step, except for one challenge step, selected by the adversary, where $\mathcal{A}$ sends a pair of records. They called a mechanism $\mathcal{M}$ $(\epsilon,\delta)$-differentially private against adaptive adversaries in the continual observation model if for any such adaptive adversary $\mathcal{A}$, the views of $\mathcal{A}$—consisting of $\mathcal{A}$’s internal randomness and the transcript of messages sent and received by $\mathcal{A}$—are $(\epsilon,\delta)$-indistinguishable when the first or the second element of the pair in the challenge step is handed to $\mathcal{M}$.
Denisov, McMahan, Rush, Smith and Thakurta (denisov2022improved, ) generalize this definition by allowing the adversary to send a pair of (data or empty) records at every step such that the sequence of first records and the sequence of second records are neighbors under an arbitrary binary relation $\sim$ on sequences of data and empty records. A mechanism $\mathcal{M}$ is said to satisfy $(\epsilon,\delta)$-DP against an adaptive adversary if, for every adversary $\mathcal{A}_{\sim}$ that adaptively sends pairs of records forming $\sim$-neighbor sequences, the adversary’s views are $(\epsilon,\delta)$-indistinguishable when $\mathcal{M}$ receives the first elements of each pair and when it receives the second elements.
@@
A mechanism $\mathcal{M}$ in the continual observation model can be viewed as an IM with a single possible initial state, representing an empty dataset. The input message set of $\mathcal{M}$ consists of the empty record and all possible data records. Under this formulation, the above definition of differential privacy can be restated as follows: For every adversary $\mathcal{A}_{\sim}$, the random variables $\mathrm{View}(\mathcal{A}_{\sim},\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A}_{\sim},\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously. Again, below we will enforce the restriction on the adversary via another IPM $\mathcal{V}$.
@@
In this alternative definition, all adversary messages are pairs of data or empty records, and the sequences of first and second elements are neighbors. Recall that in the alternative definition of DP for IMs, the first adversary message must be a pair of neighboring initial datasets, while all subsequent adversary messages must be pairs of identical queries. Next, we generalize the definition of differential privacy by allowing arbitrary conditions on the sequence of pairs sent by adversaries, without restricting the content of the messages.
@@
  • Every non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ in response. Upon receiving any subsequent input message, the CM representing $\mathcal{N}$ halts the communication.
@@
  • Mechanisms such as the Sparse Vector Technique (SVT) (dwork2009complexity, ; HardtRo10, ; lyu2016understanding, ) and the continual counter (dwork2006calibrating, ; DBLP:journals/tissec/ChanSS11, ; mcmahan2022private, ; dvijotham2024efficient, ) in the streaming settings can also be represented as CMs receiving two types of messages: question queries and data-update queries. Question queries request information about the current dataset, while data-update queries consist of instructions for modifying the dataset.